If A=2xz^2i-yzj+3xz^3k then what is the value of curlA?

1
Expert's answer
2022-01-20T17:29:39-0500

"Curl~A=\begin{vmatrix}
 i & j & k \\
 \frac{d}{dx} & \frac{d}{dy} & \frac{d}{dz}\\
 2xz^2 & -yz & 3xz^3
\end{vmatrix}
\\
i(\frac{d}{dy}(3xz^3)-\frac{d}{dz}(-yz))-j(\frac{d}{dx}(3xz^3)-\frac{d}{dz}(2xz^2))+k(\frac{d}{dx}(-yz)-\frac{d}{dy}(2xz^2))\\
Curl~A=i(y)-j(3z^3-4xz)+0k\\
Curl~A=yi-(3z^3-4xz)j"